OpenStudy (anonymous):

lim tan6t/sin2t as t->0

OpenStudy (ash2326):

@uno we know \[\lim_{t\to 0} \frac{\sin t}{t}\ \\ =\lim_{t \to 0} \frac{\tan t}{t}\ \ \ \ \ =1\] here if we have t=2t or 6t then also this is valid. Can you use this to solve the question?

OpenStudy (anonymous):

How did you get sin t/t?

OpenStudy (ash2326):

@uno if you write the expansion of sin t \[\sin t= t-t^3/3!+t^5/5!.....\] so sin t/ t= \[\sin t / t=1-t^2/3!+t^4/5!\] if t=0 then \[\sin t /t =1-0=1\]